You need 350 mL of a 40% alcohol solution. On hand, you have a 35% alcohol mixture and a 60% alcohol mixture. How much of each mixture will you need to obtain the desired solution?

Answer by fcabanski(1391)

These problems involve two equations. The first equation is the raw amount. In this case call the 35% solution x and the 60% solution y, and the amount is x + y = 350.


The second equation is the amount of alcohol. The 350 mL desired solution has 40% alcohol, so it has .4 * 350 = 140mL of alcohol. The 35% solution has .35x mL of alcohol, and the 60% solution has .6y mL of alcohol. 140 = .35x + .6y


Solve the first equation for one variable in terms of the other. x = 350-y. Substitute that into the second equation, and then solve that equation for the remaining variable.


140 = .35 * (350-y) + .6y ---> 140 = 122.5 - .35y +.6y --->17.5 = .25y ---> 70 = y.


x = 350 - y = 350-70 = 280. You need 280 mL of the 35 % solution and 70 mL of the 60% solution.

Hope the solution helped. Sometimes you need more than a solution. Contact fcabanski@hotmail.com for online, private tutoring, or personalized problem solving (quick for groups of problems.)